Gamma ray radiation falls in the wavelength region of 1.00×10-16 to 1.00×10-11 meters.

What is the frequency of gamma ray radiation that has a wavelength of 1.00×10-16 m?
1
Expert's answer
2019-09-23T08:20:25-0400

"\lambda \nu=c" where "\lambda,\nu,c" are wavelength,frequency and speed of particle.

"c=3 \times 10^{8} m/sec"

"\lambda=10^{-16} m"

"\nu=\frac{c}{\lambda}=\frac{3 \times 10^8}{10^{-16}}=3 \times 10^{24} sec^{-1}"